# SVBench: Evaluation of Video Generation Models on Social Reasoning

### v2

@@ Abstract
Recent text-to-video generation models exhibit remarkable progress in visual realism, motion fidelity, and text–video alignment, yet they remain fundamentally limited in their ability to generate socially coherent behavior. Unlike humans—who effortlessly infer intentions, beliefs, emotions, and social norms from brief visual cues—current models tend to render literal scenes without capturing the underlying causal or psychological logic. To systematically evaluate this gap, we introduce the first benchmark for social reasoning in video generation. Grounded in findings from developmental and social psychology, our benchmark organizes thirty classic social cognition paradigms into seven core dimensions, including mental-state inference, goal-directed action, joint attention, social coordination, prosocial behavior, social norms, and multi-agent strategy. To operationalize these paradigms, we develop a fully training-free agent-based pipeline that (i) distills the reasoning mechanism of each experiment, (ii) synthesizes diverse video-ready scenarios, (iii) enforces conceptual neutrality and difficulty control through cue-based critique, and (iv) evaluates generated videos using a high-capacity VLM judge across five interpretable dimensions of social reasoning. Using this framework, we conduct the first large-scale study across seven state-of-the-art video generation systems. Our results reveal substantial performance gaps: while modern models excel in surface-level plausibility, they systematically fail in intention recognition, belief reasoning, joint attention, and prosocial inference.
@@ 1 Introduction
Recent advances in video generation [23, 38, 15] have dramatically improved visual fidelity, temporal consistency, and text–video alignment. Modern diffusion- and transformer-based architectures can now synthesize dynamic scenes with striking realism, reproducing nuanced lighting, complex motion patterns, and multi-agent interactions across a wide range of environments. Yet, despite these impressive perceptual capabilities, today’s video generation models remain fundamentally socially unaware. While multi-source data fusion has been successfully applied to early warning systems [21], current video models struggle to represent why people act the way they do. They can reproduce what happens in the physical world, but struggle to represent why people act the way they do—failing to capture the latent beliefs, intentions, emotions, and norms that structure real human interactions.
@@ 2.1 Evaluation Benchmarks for Video Generation
Multimodal technologies [40, 39, 31] have progressed rapidly in recent years, laying the foundation for modeling complex signals. Early evaluations of video generation models focused on perceptual fidelity and temporal stability, adapting video quality metrics FVD[51] and human preference scores[46]. Recent benchmarks have shifted toward interpretable, axis-wise diagnostics that decompose generation quality into orthogonal dimensions. VBench[19, 20] pioneered this approach by evaluating models across fine-grained aspects—including subject consistency, background consistency, temporal flickering, motion smoothness, and aesthetic quality—using automated evaluators validated against human judgments. Ensuring stability in these technical features is paramount; for instance, the fusion of CNNs for spatial features and GRUs for temporal dependencies, optimized via genetic algorithms, has demonstrated superior stability in capturing complex sequence dynamics [22]. EvalCrafter provides a unified toolkit consolidating visual quality, content consistency, motion realism, and text-video alignment metrics for reproducible large-scale comparison. Building on this foundation, VBench-2.0[59] introduces higher-order evaluation axes such as human fidelity, controllability, creativity, physics plausibility, and commonsense reasoning, combining general-purpose vision-language models with specialized detectors. Physics-centered suites explicitly test adherence to physical laws: PhyCoBench[5] measures physical coherence with optical-flow–guided frame prediction and automated scoring aligned to human assessment; Morpheus[58] uses real physical experiments and conservation-law–based probes to benchmark physical reasoning in generated videos. While intrinsic and physics-focused benchmarks push evaluation beyond appearance, social reasoning—involving multi-agent interactions, roles, norms, intentions, and commonsense social dynamics—remains under-explored. Our work fills this gap by designing a benchmark that isolates and diagnoses social reasoning capabilities in text-to-video generation, orthogonal to (and complementary with) quality-, alignment-, and physics-oriented axes.
@@ 3.1 Seed Suite and Feasible Subset
The distribution of short-video–feasible and long-horizon tasks across the seven social reasoning dimensions is summarized in Table 1. We acknowledge that social behaviors often follow a long-tailed distribution, where rare but critical interactions are difficult to model. Recent advances in curiosity-driven cooperation have shown promise in addressing these long-tailed multi-label learning challenges [55], a principle that informs our selection of robust evaluation dimensions.

### v4

@@ Abstract
Recent text-to-video generation models have made remarkable progress in visual realism, motion fidelity, and text-video alignment, yet they still struggle to produce socially coherent behavior. Unlike humans, who readily infer intentions, beliefs, emotions, and social norms from brief visual cues, current models often generate literal scenes without capturing the underlying causal and psychological dynamics. To systematically assess this limitation, we introduce the first benchmark for social reasoning in video generation. Grounded in developmental and social psychology, the benchmark covers thirty classic social cognition paradigms spanning seven core dimensions: mental-state inference, goal-directed action, joint attention, social coordination, prosocial behavior, social norms, and multi-agent strategy. To operationalize these paradigms, we build a fully training-free agent-based pipeline that distills the reasoning structure of each paradigm, synthesizes diverse video-ready scenarios, enforces conceptual neutrality and difficulty control through cue-based critique, and evaluates generated videos with a high-capacity VLM judge along five interpretable dimensions of social reasoning. Using this framework, we conduct the first large-scale evaluation of seven state-of-the-art video generation systems. Results show a clear gap between surface-level plausibility and deeper social reasoning, suggesting that current models remain limited in their ability to generate socially grounded behavior. https://github.com/Gloria2tt/SVBench-Evaluation
@@ 1 Introduction
Recent advances in video generation [21, 36, 15] have dramatically improved visual fidelity, temporal consistency, and text–video alignment. Modern diffusion- and transformer-based architectures can now synthesize dynamic scenes with striking realism, reproducing nuanced lighting, complex motion patterns, and multi-agent interactions across a wide range of environments. Yet, despite these impressive perceptual capabilities, today’s video generation models remain fundamentally socially unaware. They can reproduce what happens in the physical world, but struggle to represent why people act the way they do—failing to capture the latent beliefs, intentions, emotions, and norms that structure real human interactions.
@@ 2.1 Evaluation Benchmarks for Video Generation
Multimodal technologies [38, 37, 29] have progressed rapidly in recent years, laying the foundation for modeling complex signals. Early evaluations of video generation models focused on perceptual fidelity and temporal stability, adapting video quality metrics FVD[49] and human preference scores[44]. Recent benchmarks have shifted toward interpretable, axis-wise diagnostics that decompose generation quality into orthogonal dimensions. VBench[19, 20] pioneered this approach by evaluating models across fine-grained aspects—including subject consistency, background consistency, temporal flickering, motion smoothness, and aesthetic quality—using automated evaluators validated against human judgments. EvalCrafter provides a unified toolkit consolidating visual quality, content consistency, motion realism, and text-video alignment metrics for reproducible large-scale comparison. Building on this foundation, VBench-2.0[56] introduces higher-order evaluation axes such as human fidelity, controllability, creativity, physics plausibility, and commonsense reasoning, combining general-purpose vision-language models with specialized detectors. Physics-centered suites explicitly test adherence to physical laws: PhyCoBench[5] measures physical coherence with optical-flow–guided frame prediction and automated scoring aligned to human assessment; Morpheus[55] uses real physical experiments and conservation-law–based probes to benchmark physical reasoning in generated videos. While intrinsic and physics-focused benchmarks push evaluation beyond appearance, social reasoning—involving multi-agent interactions, roles, norms, intentions, and commonsense social dynamics—remains under-explored. Our work fills this gap by designing a benchmark that isolates and diagnoses social reasoning capabilities in text-to-video generation, orthogonal to (and complementary with) quality-, alignment-, and physics-oriented axes.
@@ 3.1 Seed Suite and Feasible Subset
The distribution of short-video–feasible and long-horizon tasks across the seven social reasoning dimensions is summarized in Table 1.